In 2000 joe was 53 years old and tom was 23 years old
If x years before 2000,joe age was 3 times of tom age . Find x

Answer by ikleyn(52800)

.
Answer. x = 8 years.

Solution

53 - x = 3*(23-x)  ====>  3x - x = 3*23 - 53  ====>  2x = 69 - 53 = 16  ====>  x = 8.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".